### v1

[…]
1 Introduction
In today’s era, many media forms, including online news platforms, television channels, forums, blogs, and podcasts, serve as conduits of information to the public. These media continuously produce and disseminate news, presenting a complex challenge for researchers aiming to understand news impact [1]. Despite these challenges, multiple studies across various disciplines have employed techniques for analyzing news for diverse purposes. These include analyzing trends in news media [2], detecting bias in news articles [3], evaluating the effect of news on the public [4], developing tools to recognize fake news [5, 6] and more. For example, Vasiliki et al. [7] estimated countries’ peace index out from the Global Data on Events, Location, and Tone (GDELT) dataset. [^1] https://www.gdeltproject.org/ To perform that evaluation, they utilized news documents that were related to events by a machine-coded approach. This study is one example of many others who utilized event-based data to combat real-world issues [8, 9, 10]. As Liang Zhao stated ”Event analytics are important in domains as different as healthcare, business, cybersphere, politics, and entertainment, influencing almost every corner of our lives” [11]. Moreover, Zhao further claimed that ”The analysis of events has thus been attracting huge attention over the past few decades” [11].
In their simplest form, event data convert natural language reports to a dataset where each entry has the form of Date, Source_Actor, Target_Actor, Event_Code [12]. Historically, many researchers conducted their studies using human-coded data, such as The Conflict and Peace Data Bank (COPDAB), which is a longitudinal computer-based library of daily international and domestic events or interactions [13] and WEIS Project dataset - which is a record of the flow of action and response between countries reflected in public facts [14]. However, these methods relied on costly human coding efforts that, after decades of research, came to an end [15]. As a result, event data started to be generated through content analysis - using pattern recognition and simple grammatical parsing methods made it possible to develop automated coding systems. An example of such a system is the Conflict and Mediation Event Observations (CAMEO), a widely used taxonomy [16]. The advancement toward automatic coding systems allowed researchers who are interested in employing event data to study international behavior and led to the development of computational methods for the analysis and prediction of various social, political, and economic studies [17, 18]. However, the automated machine-coded approaches are far from perfect as they are highly restrictive and still suffer from low accuracy, as suggested by Leetaru and Schrodt [15] - ”Small changes in the text or the parsing method could heavily affect the document’s tagged events.”
To unlock the potential of news articles for these purposes, the development of meaningful news embeddings became a central challenge in the NLP research field, as suggested by Ma et al. [19] ”In order to facilitate the performance of downstream NLP tasks, it is rather essential to find an efficient way to represent news as continuous vectors.” Over the years multiple studies from different domains utilized news embedding in their work, some notable examples of such works are media bias prediction [3], fake news detection [20], and online news recommendations [21].
Two main approaches to news embedding emerged: The first approach involves considering the complete information of articles. The prevailing methods in news embedding usually involve capturing semantic context with complex Deep Neural network architectures. For example, Baly et al. [3] used a BERT encoder to generate its news embeddings for its media bias classification task. The second approach relies on other features, such as topic modeling, NERs, and other semantical information, to approach the problem differently. News2vec [19] is an example of such an approach, in this work, news article embeddings were generated by incorporating relationship information between articles and events and using additional topological features in a graph comprised of extracted words and features from the article textual data and associated events.
However, a major limitation of news embedding approaches is the ambiguous manner in which the embedding is constructed. News embeddings that provide good performance on several Natural Language Processing (NLP) tasks that do not necessarily succeed in association articles to events [22]. Moreover, efficiency plays a major part when constructing online news embeddings, which is usually a streaming task. Most of these works [3, 21, 23, 24] rely on textual network embedding methods and complex Deep Neural Networks (DNN) which are shown to be expensive both in memory and computation [25].
In this study, we address these limitations by focusing on a specific facet and creating a vector space optimized for event-centric embeddings. To accomplish our goal, we relied on a combination of Named Entity Recognition (NER) techniques and topic modeling. This allows us to make accurate predictions while using fewer resources and providing clear results. To accomplish this, we introduce in this work a novel method for News Article Event-Based Embedding which receives as input themes, entities, and tagged events of a large set of articles, and the output is event-based embeddings on the given news article information.
Our method receives as input the following features of news articles: entities, themes, and tagged events, then it yields as an output event-based news embeddings for the given articles. Our method is constructed in a modular structure of three stages, each stage serves as an input to the next, as can be seen in the method’s overview (see Figure 1).
  1. 1.
    Articles and Entities Processing - A large dataset of articles and entities is processed by executing multiple tasks, including removing redundancies, eliminating article duplication, applying source-based filtering, and entity parsing and removal.
  2. 2.
    Entities Embedding Generation - Given the processed dataset we trained separate word embedding models on current and historical articles to create informative embeddings for entities and themes. In this stage, each entity and theme receive a vector representation based on their co-occurrences with each other.
  3. 3.
    Articles Embedding Generation - given the entity embeddings, we deploy a Smooth Inverse Frequency (SIF) as a methodology to construct document-level vectors (see Section 2). Then, these embeddings traverse Siamese Neural Networks trained to minimize the distance between articles that share common events.
Figure 1: This figure presents the method’s entire pipeline of news embedding generation.
To test and evaluate our study, we deployed our algorithm on the GDELT project, by collecting five years worth of online news article information from a predefined set of media sources across the United States - After processing the data contained more than 850,000 unique articles and more than one million events. We tested our hypothesis by measuring the proposed models’ ability to predict whether a pair of articles discuss common events in two different settings: daily test sets and monthly test sets. We measured both the Precision-Recall Area Under the Curve (PR AUC) metric and the Receiver Operating Characteristic Area Under the Curve (AUC). The results demonstrated a significant difference between our suggested semi-supervised method which relied on event-tagged historical data, and the baseline approach: On daily aggregated test datasets: our approach yields 0.572 PR-AUC, while the baseline yields 0.546 on average. On monthly aggregated test datasets: our best approach yields 0.369 PR-AUC, while the baseline yields 0.347 on average.
This innovative approach bridges the gap between semantic and event-based context representation, providing a versatile tool for a wide range of applications requiring nuanced event-centric news embeddings.
To summarize, our main contributions are:
  • We propose a novel way to learn latent features for news articles by extending the Semi-Supervised training process of a Siamese network model to consider article-events relationships.
  • We provide access to the research community for time-relevant embeddings of more than 320,00 distinct themes and entities. Time-relevant entity embeddings were generated depending on periodical corpora, that purposely overrepresent recently published news articles.
  • We suggest a new approach for evaluating the performance of news embedding methods, by pair-wise articles’ event attribution, which utilizes diverse real-world data of relationships between articles and events.
The remainder of this paper is structured as follows. The related work in this field is presented in Section 2. The proposed method is explained in detail in Section 3. Our datasets, experiments, and final results are presented in Section 4. Finally, Section 7 concludes this paper and provides the prospect of future work.
2 Related Work
This section contains the background necessary to understand the building blocks and motivation of the research. We discuss previous works in NLP Embedding, topic modeling, Named Entity Recognition, and studies on the GDELT project.
2.1 NLP Embeddings
Representing words and documents is a crucial part of most, if not all, Natural Language Processing (NLP) tasks. It has been found useful to represent them as vectors, which can be easily interpreted, used in various operations (such as addition, subtraction, distance measures, etc.), and compatible with many machine learning algorithms and strategies. Our method for constructing news vector representations is influenced by previous research in this field [26, 27, 28].
2.1.1 Word Embeddings Methods
Word embedding methods represent words as continuous vectors in a low-dimensional space that captures words’ semantic and lexical properties, as stated by Agrawal et al. [29]. The representation of words as numerical vectors has been around for more than forty years; however, only recently have neural networks been employed for this purpose. These embeddings can be obtained from the internal representations of neural network models trained on textual data, as suggested in several works [30].
In 2013, Mikolov et al. [31] suggested two iteration-based methods: The first is the Continuous-Bag-of-Words (CBOW) model, which predicts the center word from its surrounding context. The second approach is the skip-gram model, which predicts the surrounding context words given a center word. It focuses on maximizing probabilities of context words given a specific center word.
They developed approaches that are considered non-contextual models, in which each word corresponds to a single vector, regardless of how the word is used. However, recent researchers suggest contextual models in which the training modality allows more vector representations for a word based on how it is used. For example, the word ”mean” will be represented with a single vector for non-contextual models regardless of its meaning: ”average” or ”not nice” [30]. In 2018, Devlin et al. [32] work is an example of a contextual model, which introduced the Bidirectional Encoder Representations from Transformers (BERT) model. BERT is designed to pre-train deep bidirectional representations from the unlabeled text by joint conditioning in both the left and right contexts in all layers.
In this work, we relied directly on the low-rank approximation of co-occurrences of themes and entities by using the GloVe model developed by Pennington et al. [26]. In their study, Pennington et al. proposed efficiently leveraging statistical information by training only on the nonzero elements in a word-word co-occurrence matrix rather than on the entire sparse matrix or individual context windows in a large corpus. In their results, GloVe provided better results than Word2Vec, and notably with higher efficiency, as they stated: ”the efficiency with which the count-based methods capture global statistics can be advantageous”  [26]. Therefore, GloVe can easily be adapted to the use case in this work since we did not utilize the news article’s full text but considered an efficient approach to use themes and entities while disregarding their position and textual context.
2.1.2 Sentence and Document Embeddings Methods
Sentence and document embedding methods map a feature vector representing a document or text. For instance, previous works have computed phrase or sentence embeddings by composing word embeddings using operations on vectors and matrices [33]. Another example of such an approach is proposed by Phang et al. [34], as they trained BERT, a powerful pre-trained sentence encoder.
In this study, we utilized the Smooth Inverse Frequency (SIF) algorithm for a naive representation of news article vector representations [27]. SIF computes sentence embeddings as a weighted average of word vectors by deploying a weighting scheme known as Smooth Inverse Frequency, which is meant to improve sentence embedding performance. SIF aimed to capture the significance of words in a sentence by down-weighting common words and up-weighting less frequent words [27].
2.2 Topic Modeling
Understanding the core themes associated with a document collection is a fundamental task in today’s information era [35]. The purpose of topic models is to detect those ”themes,” which are the ideas and subjects that ”connect” a set of text documents. An example of a possible theme is ”police brutality.” These models are a class of unsupervised machine-learning techniques designed for this task. In this study, we define a topic model to be an unsupervised mathematical model that takes as input a set of documents $D$ and returns a set of topics $T$ that represent the content of $D$ accurately and coherently as suggested by Kherwa and Bansal [36].
A survey by Churchill and Singh [37] mentioned several studies that employed topic models for climate change literature [38], to understand the 2016 US presidential election using newspapers [39], additional works focused on developing new methods for topic modeling while evaluating their performance on news-based datasets such as 20 Newsgroups dataset [40, 41]. [^2] Link for the 20 Newsgroup dataset: http://qwone.com/~jason/20Newsgroups/
In our work, we decided to use the Global Content Analysis Measures (GCAM) system, which runs each news article monitored by GDELT through an array of content analysis tools to capture over 2,230 latent dimensions, reporting density and value scores for each. In the overview of the GDELT project, we provide further information on their approach 2.5.
2.3 Named Entity Recognition
Named Entity Recognition (NER) aims to recognize mentions of rigid designators from text belonging to predefined semantic types such as person, location, organization, etc.” [42]. Named Entity Recognition (NER) is a versatile tool that enables the extraction of information from various sources. It is an essential component of Natural Language Processing (NLP) applications such as text comprehension [43], information retrieval [44], machine translation, and knowledge base construction [45].
The most commonly used typologies consist of a few high-level classes such as Person, Organisation, and Location [46]. There is a variety of open-source code designed for NER tasks, such as spaCy [47], NLTK [48], and flair [49]. A NER system aims to learn patterns from labeled examples, which can be used to classify new sequences of tokens.
There are mainly three main approaches for performing NER: Rule-Based, Machine Learning, and Deep Learning. Rule-based systems are interpretable and do not require training data, however designing the rules can be time-consuming and requires expertise. For example, Diez et al. [50], developed a rule-based system to recognize Medieval Spanish person names from various genres of manually transcribed texts, by defining a custom entity typology of eight main types and used multiple modules to recognize names and person attributes.
Machine-learning NER systems are statistical models that are built upon machine-learning algorithms through annotated data and selected features. Ehrmann et al. claimed in their survey [51] that ”Heavily researched and applied in the 2000s, machine learning-based approaches contributed strong baselines for mainstream NER, and were rapidly adopted for NER.”
Ehrmann et al. also claimed that deep learning techniques are dominating NER developments [51]. These deep learning approaches mainly rely on sequence labeling and use BiLSTM architectures or self-attention networks to learn sentence or sequence features [51].
Previous studies have tried to stack multiple LSTMs for sequence-labeling NER As they follow the trend of stacking forward and backward LSTMs independently, the Baseline-BiLSTMCNN is only able to learn higher-level representations of past or future. Chiu et al. [52] presented an architecture that detects word and character-level features using a hybrid bidirectional LSTM and CNN. It removes the need for most feature engineering and proposes a novel method of encoding partial lexicon matches in neural networks.
In our research, we relied on the NER algorithm output deployed by the GDELT project as input. They deployed three engines for recognizing persons, organizations, and locations by identifying an array of other kinds of proper names while attempting to compile a list of all amounts of each entity expressed in each article to offer numeric context to global events. Further description of this process can be found in the GDELT project codebook. [^3] http://data.gdeltproject.org/documentation/GDELT-Global_Knowledge_Graph_Codebook-V2.1.pdf
2.4 NLP News Embeddings
Over the years, researchers developed many supervised and unsupervised methods to represent entities, articles, or news events. In 2017, Conneau et al. [53] investigated whether supervised learning can be leveraged for sentence embedding representations. They investigated the impact of the sentence encoding architecture on representational transferability, and compared convolutional, recurrent, and even simpler word composition schemes. To evaluate their research, Conneau et al. used news articles embedded in sentences and predicted a similarity score between 0 and 5.
In 2018, Setty et al. presented Event2vec [22]. They used proposed network embedding techniques for the embedding representation of news events. As events involve different classes of nodes, such as named entities, temporal information, etc, general-purpose network embeddings are agnostic to event semantics. To address this problem, they proposed biased random walks tailored to capture news events’ neighborhoods in event networks.
In 2019, Ma et al. presented News2vec [19], which represents the context of articles, events, and additional features in a graph. They focus on both the contextual relationship between news and potential connections as the network becomes more profound, with the embedded vectors containing semantic and labeled information, as well as latent connections between different news events.
Since then, most of the works utilized news embeddings as part of a technique to achieve specific tasks [20, 21, 24], rather than suggesting a generic approach for news article embedding. For example, in 2022, Mehta et al. [20] work on fake detection included the following methodology. They embedded using graph neural network articles; then they formulated inference operators that augment the graph edges by revealing unobserved interactions between its elements, such as the similarity between documents’ contents and users’ engagement patterns.
However, most of these work for news textual network embedding methods still rely on full-text and complex Deep Neural Networks, which are expensive both in memory and computation, as Xiang and Wang claimed in their event extraction study: ”How to design an efficient neural network architecture is the main challenging issue for event extraction based on deep learning” [54]. Furthermore, the works that handle the previous limitations usually fail to collect features that do not solely describe the contextual information at the document level as stated in [19] - ”The major limitation of embedding approaches discussed above is that they produce representations and features that solely describe the contextual information at the document level”. Another limitation is that in previous works in the field, the evaluation process depends on measuring several tasks that do not necessarily determine the success of associating articles with events. In the works of Ma et al. and Setty et al. [19, 22] that concentrated on the embeddings process of news, they chose for evaluation tasks of news recommendation, or stock prediction tasks, as well as the majority of works in the field which utilized news embeddings for a specific task, such as described in the following services [5, 11, 55, 23].
There are numerous challenges to performing the task of associating events with articles, First, the positive rate is very low. In our case, on average the probability of two articles sharing a common event given they were published in the same month is less than 0.05% (see Table LABEL:tab:datasets_statistics). Second, An article can be associated with multiple events, creating a state of uncertainty in the detection process, and limiting the possible solutions. To alleviate this difficulty some previous works even chose to assume that a single article can be only related to one event [56]. Third, as we discussed earlier, robust methods that rely on full textual data, transformers, or graph neural networks often require high computational and memory resources [25, 57].
2.5 GDELT Project
The GDELT Project [15] is a real-time network diagram and database of global human society for open research which monitors the world’s broadcast, print, and web news from nearly every corner of every country in over 100 languages and identifies the people, locations, organizations, counts, themes, sources, emotions, counts, quotes, and events driving our global society every second of every day, creating a free open platform for computing on the entire world.
By employing named entity and geocoding algorithms [58], The GDELT Global Knowledge Graph (GKG) compiles a list of every person, organization, company, and location from every news report, explicitly designed for the noisy and ungrammatical world that is the world’s news media. In the GDELT project, tones refer to the emotional sentiment or attitude expressed in news articles and reports. The tones are quantified on a numerical scale ranging from extremely negative to extremely positive. The tones in GDELT data are represented as numerical values, typically ranging from -100 (extremely negative) to +100 (extremely positive), with 0 being neutral. [^4] https://analysis.gdeltproject.org/module-gkg-tonetimeline.html Moreover, GDELT provides the Global Content Analysis Measures (GCAM), [^5] https://blog.gdeltproject.org/introducing-the-global-content-analysis-measures-gcam/ this system runs on each news article monitored by GDELT through an array of leading content analysis tools, and topic models to capture and quantify latent emotional and thematic signals subconsciously encoded in the world’s media.
The GDELT Project offers a comprehensive Event Database that contains records of more than 300 types of physical activities taking place around the world. These activities can range from riots and protests to peace appeals and diplomatic exchanges. The database provides the location of the activity, whether it’s a city or a mountaintop, and covers the entire planet, dating back to January 1, 1979. It is updated every 15 minutes, making it a reliable source of real-time information. Examples of some of the identified events include ”ENVIRONMENT AND NATURAL RESOURCES”, ”LABOR MARKETS” and ”CRISIS WELLBEING HEALTH”. [^6] https://blog.gdeltproject.org/gdelt-2-0-our-global-world-in-realtime/. GDELT also considered ’mentions’ - which are the unique pairings of the article and its discussed event. Multiple articles may discuss a shared event, and one article may discuss multiple events.
Using GCAM, one can assess the density of “Anxiety” speech via Linguistic Inquiry and Word Count (LIWC) [^7] https://www.liwc.app/ project, ”Positivity” via Lexicoder, [^8] https://quanteda.io/reference/data_dictionary_LSD2015.html “Smugness” via WordNet, [^9] https://wordnet.princeton.edu/ and so forth, in total, 18 content analysis systems totaling more than are deployed. The volume of the topic models deployed by GDELT, and the accessibility they provide the output per article, make it an extremely valuable resource for our analysis.
3 Methods
The main goal of this study is to develop a method for generating event-centric news embeddings. To address this goal, we present our lightweight generic method that consists of three stages, as depicted in Figure 1. The implementation of each step can vary depending on the algorithm utilized. For example, other methods could be considered for generating entity embeddings, semi-supervised learning, article embeddings, NER, topic models, etc. In this section, we present our implementation of the method and its main components.
3.1 Articles and Entities Processing
Given a large corpus of articles, we extract and process the given articles into structured features, such as entities (persons, locations, and organizations), themes, and associated events. To accomplish this, we perform multiple steps: (a) removing redundant articles and rare entities, and handling article duplication, (b) applying source-based filtering, and (c) normalizing entities. In the following subsection, we will provide details on each step.
3.1.1 Entities and Themes Extraction
Given a corpus of textual articles $A$, we represent each article as a set of themes and entities. First, we suggest extracting those features from the textual corpus by deploying NER algorithms that extract locations, persons, dates, organizations, and the number of times each entity appeared in a given article. Additionally, we utilize topic models to extract topics and tones conveyed in the article.
3.1.2 Processing Articles Entities
The next step is to formalize the entities in each article with a global representation. Since multiple NER and topic models can be chosen, there might be some variance in how an entity is recognized. We created mappers for each entity type that maps (Entity, Entity Name), e.g., (Person, Donald Trump), then we matched similar entities to their global representation, e.g., (Donald Trump, DONALD TRUMP).
We have implemented a naive method for performing entity matching. Namely, given the entity type, a set of normalizing operations was performed. We cast dates in a unified format. [^10] We cast dates to the following formats: Day/Month/Year (dd/mm/yyyy) or Day/Month (dd/mm). We cast persons and organizations to lower cases and locations to match their country and city codes. Next, we aggregated mappers for each entity type $<Entity,Count>$ during the pipeline. Lastly, when an entity that has not appeared yet is given as input, we check for the closest entity by measuring the hamming distance with a predefined threshold. Hamming distance is a popular measure of the distance or similarity between objects. The hamming distance is defined as the number of positions at which the corresponding symbols differ [59]. If the new entity does not have a match, it is inserted as a new entity to the relevant mapper.
In many cases, large datasets of entities contain frequency distributed in a power-low manner (see, for example, Figure 2), meaning a few entities appear in most of the article’s data. To optimize runtime and memory utilization, we handle the vast size of the entities by removing the entities that appeared less than the selected thresholds in the cleaned dataset. We have manually defined the threshold parameters for each entity type based on the given data and the loss percentage of unique entities.
Figure 2: This figure contained a Pareto chart of the Top 35 persons occurrences in the full collected dataset from GDELT.
3.1.3 Cleaning the Data
After processing the entities and themes, the next step is to clean redundant data. We consider two types of redundant articles: duplicated or remarkably similar and very short. There are various reasons for article duplication. For example, the same article from the same source could be parsed again in the given dataset, or the same article was obtained from a different source. We decided to remove articles with the same themes, entities, and tones (provided with GCAM), as articles we have already processed to mitigate over-representation in data of the same popular articles.
To handle the redundant short article, we removed all the articles containing less than a given predefined threshold of valid entities and themes.
3.1.4 Splitting the Data
Lastly, we deployed a splitting mechanism after completing the previous processing tasks. Given a time frame, it will split the given dataset of articles into smaller datasets, such as daily and monthly separated datasets. We perform this for several reasons: First, the embedding representation of entities and articles should based on the relevant reality - for example, currently, we expect Joe Biden’s embedding representation to be closer to the US President entity than Donald Trump’s embedding. Moreover, to keep the method lightweight, we consider only part of the dataset at all times if necessary. During this study, we noticed that most events have a short life period. Therefore, the information loss can be contained with the predefined time-frame split.
3.2 Embedding Generation for Entities
This stage of the method is responsible for producing embedding representations for entities. Given time-separated datasets of processed entities found in articles, this stage generates embedding for each entity. To perform this, we train GloVe models for each time-partitioned dataset, based on the entities co-occurrences in the articles. To improve the generation of time-sensitive embeddings for entities, we developed a pooling mechanism of additional articles before the training process to represent current and historical information.
3.2.1 Embedding Generation for Entities - Adapting GloVe
To generate entity embeddings, we utilized the GloVe algorithm [26]. We adopt the GloVe algorithm in the following manner: Instead of entering a list of words that appeared in a sentence, the GloVe model receives the set of entities and themes that appeared in the same article as input. We defined the context window size as the length of the largest set of entities in all the articles. Since we do not have the order of the entities that appeared in the articles, there is no meaning for a limited context window. The GloVe model incorporates global and nearby entities’ data and attempts to assimilate the benefits of the neural network language model.
We use the following notations:
$X_{ij}$ is the number of times entity $i$ occurs in the same context window (article) as entity $j$;
$X_{i}=\sum_{j}X_{ij}$;
$P_{ij}=\frac{X_{ij}}{X_{i}}$; is the probability that entity $j$ appears in the same article as the entity $i$.
$W_{i}$ is a vector embedding representation of entity $i$. For each entity, we define the following equation:
$log(X_{ij})=w^{T}_{i}w_{j}+b_{i}+b_{j},$
where $b_{i}$ and $b_{j}$ are the biases of the central word and the context respectively.
Overall, the loss function for GloVe preparation is displayed in the accompanying condition:
$J=\sum_{i,j}f(X_{ij})(w^{T}_{i}w_{j}+b_{i}+b_{j}-log(X_{ij}))^{2}$
Where $f(X_{ij})$ is a weight function.
Before training the GloVe model, we developed a pooling mechanism to ensure that the entities’ embeddings generated in a specific period contain recent and historical information. The pooling mechanism was constructed as follows: First, we sampled articles from previous periods in a decreasing manner, adding them to the current period corpus used to train the model. It is important to note that we did not sample data from the future to prevent leakage or bias to the entities’ embeddings. To define the pooling mechanism we use the following notations: The dataset of all the processed articles’ entities that appeared in a set time frame $t$ is marked as $M_{t}$. $rand(M_{k},p)$ is a function that randomly selects articles’ entities without replacement with $p$ probability to pick an article. The article corpus for each time frame was assembled in the following manner:
$Corpus_{M_{k}}=\{{M_{k}},rand(M_{i-1},p_{1}),...,rand(M_{n},p_{n-1})\}$
3.3 Embedding Generation for Articles
We utilized two distinct techniques to create news article embeddings. Firstly, we deploy a sentence embedding model on the embedded entities space so that given a set of embedding representations of article entities, this model generates a vector representing the article. Secondly, we train a semi-supervised DNN, similar to Ohi et al. work [60]; we did so on the tagged article event data. This model receives input from the naive articles embedding to produce a new vector space that considers the cross-document relationships given shared events. After completing the two approaches, we generate a third and final representation for the article embeddings - where we rely on the synergy of the two previous methods.
In the following sections, we describe the steps required to generate our method’s final article embedding: deploy sentence embedding model (see Section 3.3.1), then train a DNN semi-supervised model (see Section 3.3.2). Lastly, we perform a synergy based on the previous steps (see Section 3.3.3).
3.3.1 Embedding Generation for Articles Using SIF
In this stage, we construct a naive article embedding representations using the SIF algorithm [27]. Conceptually, to represent a complex object one can apply a simple aggregator function on the object’s component embeddings. By this logic, the SIF model receives as input a list of the entities’ embeddings that appeared in each article, as generated in the previous step (see Section 3.2), and outputs article embeddings.
We chose the SIF model as the aggregator function [27] due to its simplicity and competitive performance. The underlying intuition is that the embeddings of too frequent words should be down-weighted when summed with those of less frequent ones is compatible with the use-case of entities in articles - frequent entities provide less context required for distinguishing articles from one another.
We implemented this and considered it as the base approach for the task of article embeddings. However, there might be a drawback to such an approach, it might suffer from underfitting. The simplicity of the SIF model does not account for additional information given by a group of certain entities. To uncover that pattern one can use the information between articles - in our case articles that discuss common events.
Figure 3: The architecture and training process of each Triplet Siamese Network.
3.3.2 Embedding Generation for Articles Using DNN Semi Supervised
We chose the Siamese Network model for this task, introduced by Bromley et al. [61], based on two identical Deep Neural Networks. The central concept of the approach is independent of the complexity of the model. Therefore, we chose a basic architecture for the models, which is independent of the specific time frame of the training data. This means that we applied the same model architecture to the entire dataset. The input layer of the model equals the size of the embedding generated by the SIF model. Three hidden layers are densely connected. The output layer of the model equals the input embedding size. The ReLU activation functions were configured on top of each layer, with uniform kernel initialize functions and L2 regularization functions. Figure 3 presents an overview of the suggested implementation of the Siamese Network.
Similar to Ren et al. [62], we have also trained a Siamese Network model, with a Triplet Loss objective. Therefore, in our implementation, the architecture of the Siamese model contained three identical DNNs, each of which was built based on the architecture properties we described above. Triplet Loss objective is a distance-based loss function that operates on three inputs:
  1. 1.
    anchor (a) is any arbitrary data point.
  2. 2.
    positive (p) which is the same class as the anchor.
  3. 3.
    negative (n) which is a different class from the anchor.
Namely, it is defined as:
$L=\max(dist(a,p)-dist(a,n)+\epsilon,0).$
Our aim in using such network architecture is to generate article embeddings. For our problem, we define the anchor as an article and we consider an article that shares common events for the anchor as positive and an article that does not consider a negative sample respectively.
We deployed the following triplet mining policy: To construct a triplet for a particular anchor article $a$, we must select a positive article, $p$, that shares at least one mentioned event, and a negative article, $n$, that does not share any common event with $a$. In a dataset with $N$ training articles, there are $O(N^{3})$ possible triplets, many of which do not help the training converge (e.g., triplets where $dist(a,n)>>dist(a,p))$.
Many works have discussed the benefit of hard negative mining in constructing triplets that produce useful gradients and therefore help triplet loss networks converge quickly [63, 64]. However, previous works also show that hard positive examples can increase clustering within a class [65]. Therefore, policy within each batch and optimized our loss function using ADAM optimizer [66] - we first sample $N$ triplet pairs, and out of those we select the top $K_{1}$ most violated matches in each sample. This is done by computing the pairwise distances between all articles within the initial batch. The pairwise distances were calculated as follows:
$dist(a,p)-dist(a,n)=\sum(a-p)^{2}-\sum(a-n)^{2}.$
To avoid clustering within a specific class, in addition to the $K_{1}$ most violated matches, we added to the batch $K_{2}$ random matches from the initial sample.
As suggested by Roy et al. [67], to generate the article embedding using the trained Siamese model, we consider a single DNN from the Siamese network triplet and utilize it to produce the article embeddings. We do so by taking the output of the last layer in the selected DNN, that is the same size as the input layer.
3.3.3 Embedding Generation for Articles - Synergy
After the completion of the training process of the Siamese network, we generated two different embedding representations for all articles given their entities: (a) the naive SIF approach, and (b) the data-driven Siamese network approach.
This stage relied on the premise that there is synergy between those approaches; each approach has its merit, and depending on the data, each might contribute additional information. To consider this, in the final stage, the algorithm receives as input those embedding representations and outputs a vector of their concatenation, which represents an event-centric embedding of the given article.
3.3.4 Articles Embedding Evaluation
As explained in the previous sections, each method generated an embedding representation for each article by receiving the article’s entities embedding as input. We perform extensive experiments to evaluate the performance of the proposed methods for event-centric article embedding. We measure the methods’ performances on the following task: Given a periodic dataset of articles’ embeddings and their tagged events, we create pairwise permutations of all the articles in the given dataset. A pair of articles is considered close or instead tagged as positive if both articles share at least a single joint event. A simple Euclidean distance between each pair’s embedding representation concludes their ranking score.
To validate the performance of our method, we performed a cross-dataset evaluation on different periods across the collected data without fine-tuning the models with the variant training data. The evaluation process considered the three proposed methods for article embeddings. To measure the methods we consider the task of common event attribution, and evaluate two metrics: precision-recall AUC, and ROC AUC [68]. Moreover, we also conducted the Friedman test [69] and Nemenyi post-hoc [70] to determine whether a significant statistical difference exists between the article embedding generation methods.
4 Experimental Setup
We evaluated our methods (see Section 3 and Figure 1) on monitored news sources by GDELT from 02/2015 to 01/2021. Finally, after the embedding generation for articles is complete, cosine similarity between all the article pairs in the same time frame is computed, to determine which articles share at least a single common event. The dataset is described in detail in Section 4.1, the experiment is described in Section 4.2, and the results are presented in Section 5.
4.1 Data Description
Throughout this paper, three different types of information are employed, including global knowledge graphs, events, and article-event mentions. These datasets are accessible by GDELT as three different types of files, for additional information on the GDLET source data (see A.2). The period of the data that have been collected is 02/2015 to 01/2021. Overall, in this research, we have evaluated 850,492 unique articles, 1,002,967 events, and 3,551,906 mentions.
We decided to collect all the articles from various popular media sources; each is viewed significantly more by people who identify as liberals or conservatives. We depended on previous works that classified different sources to a specific political view, such as described in a study by Iyengar et al. [71]. We have also decided to monitor the affiliates of those networks if they exist. For example, Fox News has more than 50 local affiliates, so we collected data on the articles from all of Fox News affiliates and classified them as the same source ’FoxFamily’. Including affiliates, we have gathered articles from 605 media sources (see A.2.1). Figure 4 presents the distribution of media sources in the collected data sets, labeled by political affiliation.
Figure 4: This figure maps each monitored media source to the articles published by it in the preprocessed dataset. Conservative and Liberal media sources are labeled in blue and red respectively.
Table 1: Yearly aggregated datasets statistics used in this work for performance analysis.
Year #Articles #Events #Mentions %Monthly PR %Daily PR
2015 78877 96783 361148 0.0547 0.5012
2016 166754 196716 702328 0.0423 0.3630
2017 162264 199470 694312 0.0436 0.3764
2018 160505 192327 681872 0.0470 0.4066
2019 139823 157397 555292 0.0398 0.3515
2020 142269 160274 556954 0.0381 0.3253
- 850,492 1,002,967 3,551,906 0.043% 0.377%
4.1.1 Labled Datasets Construction
We validate the proposed methods by performing cross-dataset testing by using our method. We have separated the entire collected article data into 66 datasets separated by months. We derived from the GDELT open-sourced data, Table LABEL:tab:datasets_statistics presents the characteristics of each monthly dataset included in the experiment. While Table 1 shows yearly aggregated statistical information of those datasets. It is important to note that some datasets were relatively smaller in the number of articles, events, and mentions of appearances. The difference in the datasets’ sizes led us to assume there was an inconsistency in the raw GDELT data.
We evaluated our method using these datasets on two tasks: monthly and daily attribution of two articles to share events.
  • Monthly Mentioned Events: A large-scale dataset consists of every possible permutation of 2 articles in the same month. Given the embedding representation of each pair of articles, we calculate a prediction probability score that those articles shared at least a single joint event. It is worth noting that working with this dataset is challenging because it is highly imbalanced, and the probability of pairing articles with relatively far creation dates is substantially lower since events’ mentions appearances over time were skewed toward their creation date.
  • Daily Mentioned Events: A smaller-scale dataset consists of every possible permutation of 2 articles published on the same day. Each pair of articles is assigned a probability score for that pair of articles to share a common event. It is worth noting that this dataset is less challenging because it is considerably more balanced than the monthly use case.
4.2 Experiments
In this section, we provide the implementation details and discuss the evaluation process of each step of the proposed methodology. For additional implementation notes, regarding the research experiments, see Section A.1.
Articles and Entities Processing. First, as we mentioned in Section 2.5, GDELT publishes the output of its methods to extract Persons, Locations, Organizations, Dates, Themes, and Tones. So, we extract those from the GDELT raw data and construct entries that contain the article: URL, entities, themes, events, time, and tones.
In the processing of the entities task, we define the following parameters:
  • The entities’ pruning thresholds were set as (Themes, 20), (Persons, 75), (Organizations, 50), (Dates, 25). Locations entities were not removed, as they contain relatively a much smaller set of unique values. Those parameters were manually set with the condition of keeping the most common 33% of the entities for each entity type.
  • For determining whether an entity already appeared, the hamming distance similarity threshold was set to two. We have picked the threshold by manually selecting a few entities and searching their closest entities by hamming distance. We noticed that for the majority of cases, a hamming distance of two or less was caused by a typographical error.
  • To handle the redundant short articles, we removed all the articles containing less than four valid entities and themes. We assume that an article with such a few number of entities is irrelevant or there was an error in its entities extraction process.
Figure 5: Siamese-Network: Train Triplet Loss. The X-axis is the monitored training steps, every 4 steps the average loss was calculated. The labels represent each Siamese model that trained for a given month.
Embedding Generation for Entities. We decided to collect a corpus of the recent six months, using the pooling mechanism we defined in Section 3.2.1.
The GloVe parameters were set as learning rate = 0.05, max loss = 10, and vector size = 175. We fixed these parameters for all the experiments.
We manually evaluated the generated GloVe embedding by doing the following: We selected 50 different themes and entities, overall. For each entity or theme, we searched for the five closest neighbors on a monthly trained GloVe model. We had two criteria for selecting each entity:
(a) The GloVe model successfully accounted for recent information regarding that entity. For instance, we checked the closest entities to Nancy Pelosi’s embeddings during January 2021, the month of the storming of the Capitol, in which protesters took over Nancy Pelosi’s office. We checked if the embedding of the ’White House’ was a close neighbor.
(b) The embedding of that entity accounted for historical information. For example, we used the entity ’9.11’ and checked if we found the theme ’TERROR’ in its close neighbors.
After manually examining the selected entities and themes, we considered their relationships with others to be reasonable. For examples of the effect the pooling mechanism had on the entities embeddings (see Section A.3).
Embedding Generation for Articles. In each experiment, we split the events in the given dataset by a 70%-30% split to train and test by their order of appearance, as suggested by Nguyen et al. [72]. Nguyen et al. stated that ”The 70–30 split, widely employed in machine learning, strikes a balance between an adequate training set size and a sufficiently large test set, ensuring a robust evaluation of the model’s performance” [72]. In addition, data leakage articles duplication between the train and test set were removed.
We conducted this on each monthly dataset and evaluated suggested methods and the baseline. When training the Siamese Networks we defined fixed values for the hyper-parameters: ADAM optimizer with a Learning Rate of 0.0015. The Siamese Triplet model contained three DNN network architectures simple input and output layers of 175 units with three hidden layers of 128 units, we have also defined L2 regularization on top of each layer of the model. Each batch the size of 64, which consisted of 50 random pairings of articles and 14 hard pairings, as described by our defined triplets mining policy. [^11] The models were not trained on epochs but on 10,000 batches because of the sheer volume of the pair article’s permutations. We chose 10,000 batches since we saw a convergence across the datasets beforehand, which made sense in a relatively small network as we constructed. This can be exemplified by the train loss of the Siamese models we train in 2020 monthly datasets (see Figure 5). Each label is the relevant Siamese model in a given month. We constructed the embeddings, as defined in 3.3.2.
Performance Evaluation Figure 5 presents the training triplet loss curve of the Siamese network to visualize the training process on each dataset in 2020. We observed that the training loss of the models converges to a similar loss value.
5 Results
In this section, we present the results of the experiments as described in the preceding section 4.2.
5.1 Periodical Theme and Entity Embeddings
In this section, we present the number and type of each periodic embedding we generated in this work. As part of our study’s methodology, we have developed a new way of generating embeddings that are relevant to different periods. Specifically, we have created monthly embeddings for every valid entity or theme that we have identified as defined in Section 3.1. This means that even if a valid entity appears only once in all of the monitored media sources, in a given month, we can still create a monthly relevant embedding representation for it by using the GloVe model training process, as we described in Section 3.2. Our methodology involved monitoring various features from the news articles, including Person, Location, Date, Organization, and Theme. We constructed monthly embeddings for distinct entities from 02/2015-01/2021, and the count of these entities is presented in Table 2. Throughout the 66 months of monitoring, we aimed to generate 331,630 embeddings for news entities and themes each month.
Table 2: This table presents the number of unique generated embeddings per monitored entity type.
Entity Type Count
Person 202,006
Organization 103,201
Theme 16,577
Date 9,552
Location 294
Total 331,630
Table 3: This Table contains the mean and standard deviation, of each method on both daily and monthly common event attribution tasks, for PR and ROC AUC metrics.
[…]
max width= Method Monthly Evaluation Daily Evaluation Precision-Recall AUC ROC AUC Precision-Recall AUC ROC AUC SIF $0.348\pm 0.048$ $0.958\pm 0.015$ $0.547\pm 0.055$ $0.955\pm 0.014$ Siamese Network $0.344\pm 0.054$ $\mathbf{0.974\pm 0.012}$ $0.548\pm 0.060$ $\mathbf{0.972\pm 0012}$ Concatenation $\mathbf{0.369\pm 0.048}$ $\mathbf{0.974\pm 0.010}$ $\mathbf{0.572\pm 0.055}$ $\mathbf{0.972\pm 0.010}$
(a) Monthly Precision-Recall AUC Comparison
(b) Monthly ROC AUC Comparison
(c) Daily Precision-Recall AUC Comparison
(d) Daily ROC AUC Comparison
Figure 6: Statistical Analysis of Article Generation Methods - Comparison of performance using the Friedman and Nemenyi tests with a 5% significance level in the daily and monthly event attribution task.
5.2 Common Event Attribution Tasks
As described in section 4, we ran the experiments on all 66 distinct monthly datasets we collected and processed from the GDELT project. The results of those experiments appear in Figures 7. To summarize these results, we also constructed the Following Table 3 which contains the average performance of all methods, across all datasets, measured metrics, and tasks. In Figure 7, we present the performance of each method on both tasks: daily and monthly common event attribution across all 66 monthly separated datasets of GDELT data. This plot enables further investigation of the benefits and shortcomings of each method, by presenting their relationship and behavior given a dataset.
Additionally, in Figures 6, we visualize the distribution of the performance of all methods in the context of metrics: ROC and PR AUC, and tasks: monthly and daily, across all the datasets. Moreover, we performed statistical analysis to determine whether there was a significant difference in the methods. We concluded that the concatenated approach provides better performance in both tasks compared to the standalone baseline and SiameseNet by the Precision-Recall AUC evaluation.
(a) Precision-Recall AUC Comparison
(b) ROC AUC Comparison
Figure 7: ROC and Precision-Recall AUC results of the daily and monthly common event attribution task across all methods and datasets.
5.3 Statistical significance evaluation
We perform statistical analysis to verify the existence of significant differences among the evaluated methods. As suggested by [73], we use the Friedman test [69] on the results of the aforementioned approaches to identify which obtained better results regarding each designed task and metric - across all datasets.
The tests were used with a significance level of 5%. Specifically, the rank results established by the statistical the method provides the criterion for evaluating the approaches, considering that the higher the value of the calculated rank, the better the result will be in this approach. In our case, the null hypnosis was rejected in each context, as can be seen in Table 5. The rejection of the null hypothesis of the Friedman test indicates the existence of statistical differences between the methods but does not define which of them are statistically different. For that, the Nemenyi post-hoc test [70] was applied to make multiple comparisons using the values of the percentages of means for each approach evaluated in all the executed datasets in relation to their context, as can be seen in Table 6. The results of the statistical analysis are presented in A.4.
Therefore, in all cases, there were significant differences between the baseline SIF approach and the Concatenated approach. While by the Precision-Recall metric, there have also been significant differences between our proposed SiameseNet and the Concatenated approach.
6 Discussion
Our results point to the following important notes:
First, the utilization of entities, themes, and events to construct news embeddings has shown promising results, this is reflected in Tables 3, 1. Even though the task of shared event detection between pairs of articles in a said timeframe is challenging as discussed in the Introduction section, all proposed methods achieved extremely high ROC AUC scores, all above 0.95. While also providing impressive results with regard to the more challenging metric of PR AUC.
Second, the utilization of the semi-supervised approach, which considers also cross-document latent information, given the tagged events appears to contribute to our method performance. From Figure 6 it is evident that the semi-supervised approach achieves significant improvement compared to the SIF vaseline by the ROC AUC metric on all tasks. Additionally from analyzing Figure 7, we see that even though there is a high correlation between the method’s performance, a variance in performance exists. Meaning we found some cases in which the baseline performs better than the semi-supervised approach. We attribute this behavior to two reasons: (a) we found a moderate correlation of 0.369 for each dataset size and the gap in performance between the SIF and semi-supervised approach, more data led to a bigger gap in favor of the Siamese network, (b) a lack of adaptation of the Siamese network to the given dataset or even task led to variances in its results.
Third, our method which generates news article embeddings was able to produce embeddings for both articles and entities in the process, on a huge scale of news data, without the use of GPU in the process. Using the resources described in A, we ran every step of the pipeline, and every experiment on all the generated datasets in parallel. For each entity and theme we monitored, over 320,000 overall (see Table 2), we generated topical embeddings for every period they appeared in from 07/2015 to 01/2021. We have considered all articles from dozens of different monitored media sources and their affiliates, as presented in Figure 4, resulting in a dataset of over 850,000 articles. This enables researchers with fewer resources to investigate the large-scale of online news, using this method.
Fourth, we constructed topical monthly entity embeddings for over 320,000 distinct entities and themes. The pooling mechanism that controlled the training set of the GloVe algorithm served as a critical point in this research. This mechanism not only allowed us to control the resource utilization of our suggested method but also fine-tuned the representation of each entity to a recent time frame as its articles’ publication date. In Table 4 we further explain and demonstrate real-world examples of the importance of time topical embeddings. We plan to share these topical entity embeddings with the research community. These topical embeddings could serve other studies that seek to examine how relationships between entities change over time.
Fifth, The experimental results, as presented in the previous section, demonstrate the superiority of our proposed method compared to the baseline method SIF and the Siamese network. This can be detected in both Figures 6, 7, and in Table 3. We detected only three instances out of 66 datasets and two tasks in which the SIF approach surpassed our technique. Across both daily and monthly event attribution tasks, our technique consistently outperformed SIF, surpassing it by 2.57% in PR-AUC and 1.78% in ROC-AUC for the daily task, and by 2.15% in PR-AUC and 1.67% in ROC-AUC for the monthly task. We did see, that in the less restrictive metric of ROC AUC, the semi-supervised approach and the concatenation approach received similar results. However, as brought from the statistical analysis, performed using the Friedman and Nemenyi tests with a 5% we validated the significant improvement of the concatenation approach on the PR AUC metric, on both tasks. As detailed in Section 5.3 and A.4.
Lastly, we observed that in the cases in which a better performance is yielded by the Siamese network or SIF Methods compared to the Concatenated approach, there was a high gap between the SIF approach and the Siamese approach, leading to failure of the concatenated approach to successfully account for additional information by utilizing both approaches. Resulting in lesser performance than taking the best approach given the dataset.
Finally, the methods and method presented in this research have some limitations worth noting: First, it is reliant on the historical existence of event tagging for news articles. Which might not be available, or reliable enough. Second, as we suggested by the results our methodology is rather static and the pipeline should preferably be manually tuned to provide adaptive embeddings to the required task, these concerns both in the semi-supervised training process and in the feature selection phase.
7 Conclusions and Future Work
This research has presented a novel approach for generating event-based embeddings for news articles, addressing the limitations of existing methods that rely solely on the textual content of documents, disregard runtime and memory costs, and do not rely on cross-document information. By leveraging the historical context of entities, themes, and their associations with specific events, our proposed method aimed to capture the latent relationships between these elements, ultimately enhancing the representation of news articles in a vector space.
We showed our method’s ability to produce articles’ embeddings that consider the article’s published time, represent the latent event information that appears in them, and do so in relatively low resource usage on large-scale datasets. We also demonstrated the method’s ability to surpass the arbitrary aggregation function of the baseline SIF approach. Moreover, we utilized the synthesis between the SIF approach and the semi-supervised Siamese network approach.
This research offers multiple directions for possible future work:
investigating the interpretability and explainability of the generated embeddings could be a valuable research direction. Understanding the factors that contribute to the embeddings’ representations and their relationships with events and entities could lead to more transparent and trustworthy models, as well as facilitate the incorporation of human expertise and domain knowledge into the embedding generation process.
This also paves the way for analyzing the relationships between certain entities and tense or biased media coverage. In the future, we may apply a holistic approach to event clustering based on the methods proposed in this work, to facilitate possible research in the media slant or political news.
Another interesting opportunity for research could be to suggest an automated methodology for training a flexible semi-supervised network that might result in substantial improvement in the results, given the dataset it receives as input. This makes sense, considering the high variance of both dataset size and the positive rate as presented in depth in Table LABEL:tab:datasets_statistics.
Acknowledgments
We would also like to thank Dr. Kalev Leetaru the Founder of the GDELT Project, for his assistance with understanding the GDLET project, and the possibilities its data contained.
In addition, while drafting this article, we used ChatGPT for editing according to necessity. We also used Grammarly to check the grammar, spelling, and clarity of the text in this study.
Appendix A Research Notes
In the appendix, we present the detailed implementation notes of our method, including the running environment and the parameter settings for the deployed algorithms and Neural Networks. Finally, we provide further references to the GDELT raw data.
A.1 Implementation Notes
Running Environment. The experiments are conducted on a single Linux server with AMD EPYC 7702P 64-Core Processor and 512Gb RAM, no GPU used for train or evaluation. Our methods are implemented on Tensorflow 2.4.0, Keras 2.7.0, and Python 3.8.5.
Hyperparameter settings. Adam optimizer with a Learning Rate of 0.0015, The network architecture was a simple input and output layer of 175 units with three hidden layers of 128 units, we’ve also defined l2 regularization on top of each layer of the model. The models were trained on 10,000 epochs, and a batch size of 64, which consisted of 50 random pairings of articles and 14 hard pairings, as described by our defined triplets mining policy. The architecture of the Siamese network is presented in Figure 8.
Figure 8: This figure presents a summary of the proposed SiameseNet Architecture.
A.2 Data Notes
For further references on the GDELT raw data see the GDELT event database format codebook  [^12] http://data.gdeltproject.org/documentation/GDELT-Event_Codebook-V2.0.pdf, and also GDELT Global Knowledge Graph (GKG) data format codebook  [^13] http://data.gdeltproject.org/documentation/GDELT-Global_Knowledge_Graph_Codebook-V2.1.pdf.
To provide a better indication of the pitfalls of news event association, we will provide real-world examples of articles that have been published, a short NER output on a sample of their content, and their event association by GDELT:
Figure 9: These images provide examples of classic NER output on two US 2020 Election Debate News Articles
Figure 10: These images provide examples of classic NER output on two Iran Meddling with US 2020 Election News Articles
Examining the two articles in Figure 9, and the NER output on their content it seems fairly obvious that those discussed a common event: the upcoming second presidential debate. The logic of utilizing entities and themes to detect events is crystal clear here. As expected, the GDELT automated coding system agrees with our analysis, only it associated the NBC article with another event, that unfortunately is relevant only to this specific article in all of our collected datasets. On that note, one could claim that the Fox News article discussed another event: the White House 2012 Christmas Pary. This represents the problematic nature of event association and the difficulty behind it.
Another real-world example is presented in Figure 10, again the NER output of the articles’ content marked in it. Just like in the previous Figure 9, it seems that this pair of articles is extremely similar and therefore shares an event, we would label this event as something like ’Iran Meddling with 2020 US Election’. However, by GDELT annotations those articles share no common event, even though they share similar context, entities, and themes.
A.2.1 Source Data Description
The following are the descriptions of the files generated by GDELT and used in this work, to construct the datasets.
  • GKG Files: These files contained information on each document, such as an identifier (URL), date of the article, source, and most importantly its entities, the entities we collected throughout this paper were - Themes, Persons, Organizations, Locations, and Dates.
  • Mentions Files: These files connect between events and articles. They contain information on the relationships between them, such as the global event identifier, event time date, article time date, article identifier, the actors of the event, and so forth. They also contain the certainty score in the connection. We considered only mentions with absolute certainty by GDELT.
  • Export Files: These files contain information on the events. They contain information on the relationships between them, such as the global event identifier, event time date, article time date, article identifier, the actors of the event, and so forth.
Table LABEL:tab:datasets_statistics provides a comprehensive overview of the attributes included in each monthly dataset that has been generated, from the GDELT source data, in the period of 01/2015 - 01/2022.
A.3 Embedding Generation Notes
In this section, we will provide examples of the effect our methodology had on the embedding generation process.
Table 4: This table, contains the closest neighbors of a given entity, by a trained GloVe model on month 01/2016, with and without the pooling mechanism.
Entity Entity Neighbour GloVe GloVe+Pool Mechanism
9.11 First ECON_SOVEREIGN_DEBT 11.9.2001
Second Marlon_Brando World_Trade_Center
Third Stockard_Channing Memorial_Museum
Fourth Elizabeth_Taylor WB_2467_TERRORISM
Fifth Michael_Jackson TAX_FNCACT_FIREMEN
Vladimir_Putin First TAX_ETHNICITY_RUSSIANS TAX_ETHNICITY_RUSSIAN
Second TAX_ETHNICITY_RUSSIAN TAX_WORLDLANGUAGES_RUSSIAN
Third TAX_WORLDLANGUAGES_RUSSIAN RS
Fourth Alexander_Litvinenko TAX_ETHNICITY_RUSSIANS
Fith RS Alexander_Litvinenko
We will investigate how the pooling mechanism affected the generation process of entity embeddings 3.2. An example of the pooling effect is exemplified by the examples in Table 4. The Table showcases two examples of how data reliant is the Glove model. The single monthly trained GloVe model was not able to represent the 9.11 entity. It was, however able to successfully encapsulate the Vladimir_Putin entity, while including Alexander_Litvinenko as a close neighbor.
Alexander Litvinenko was a former officer of the Russian Federal Security Service (FSB) who defected and became a British citizen. He died suddenly from poisoning in 2006, leaving many questions unanswered. However, in January 2016, a public inquiry concluded that the two suspects who carried out Litvinenko’s murder were likely acting under the direction of the FSB and with the approval of Russian President Vladimir Putin. The GloVe, both with and without the pooling mechanism, captured the topical relationship between Litvinenko and Putin during that period.
A.4 Statistical Evaluation Notes
To perform the statistical evaluation Friedman test was used, the test evaluated whether a significant difference in the result exists between each pair of methods. The results are described in Table 5
Table 5: Friedman test results per context - task and metric
Task Metric P-value
Monthly Precision-Recall AUC 3.95E-14
Monthly ROC AUC 5.33E-21
Daily Precision-Recall AUC 3.94E-19
Daily ROC AUC 3.18E-22
Table 6: Nemenyi post-hoc test results per context - task and metric
Task Metric Method 1 Method 2 P-value
Monthly Precision-Recall AUC Concatenation SiameseNet 0.001
Concatenation SIF 0.001
SiameseNet SIF 0.8447
Monthly ROC AUC Concatenation SiameseNet 0.8447
Concatenation SIF 0.001
SiameseNet SIF 0.001
Daily Precision-Recall AUC Concatenation SiameseNet 0.001
Concatenation SIF 0.001
SiameseNet SIF 0.8945
Daily ROC AUC Concatenation SiameseNet 0.9
Concatenation SIF 0.001
SiameseNet SIF 0.001
Table 7: Datasets statistics are used in this work for performance analysis.
Dataset Statistics
Year Month #Articles #Events #Mentions %Monthly PR %Daily PR
2015 7 14961 18746 69644 0.074% 0.637%
8 12872 16743 58580 0.052% 0.496%
9 13701 16772 64938 0.065% 0.528%
10 11723 14013 51964 0.040% 0.361%
11 13389 15968 61664 0.057% 0.512%
12 12231 14541 54358 0.040% 0.473%
2016 1 12880 14809 55648 0.040% 0.340%
2 13763 16514 60482 0.042% 0.353%
3 14316 17174 60814 0.037% 0.327%
4 13285 15285 53896 0.037% 0.301%
5 14139 16923 57166 0.050% 0.402%
6 14860 18055 62852 0.051% 0.414%
7 15484 17932 71014 0.035% 0.355%
8 12916 15085 52476 0.042% 0.320%
9 15010 16837 61678 0.045% 0.369%
10 13566 16402 55620 0.036% 0.324%
11 13498 16492 57346 0.054% 0.454%
12 13037 15208 53336 0.039% 0.397%
2017 1 13989 17375 59614 0.041% 0.328%
2 12799 15761 51010 0.039% 0.304%
3 15070 18638 63242 0.042% 0.375%
4 14166 17521 61750 0.042% 0.331%
5 14880 18419 64636 0.040% 0.358%
6 13616 16399 58848 0.039% 0.364%
7 12376 14882 53130 0.055% 0.407%
8 13325 16253 55712 0.062% 0.524%
9 13458 16543 59196 0.040% 0.366%
10 13996 16977 61068 0.034% 0.305%
11 13108 16295 57012 0.041% 0.357%
12 11481 14407 49094 0.048% 0.498%
2018 1 12445 15108 51280 0.045% 0.370%
2 11923 14758 49846 0.049% 0.355%
3 15170 18369 63828 0.040% 0.365%
4 14326 17520 63204 0.047% 0.416%
5 15436 19194 68934 0.048% 0.443%
6 14167 17737 61436 0.051% 0.410%
7 13790 16397 59174 0.049% 0.408%
8 15655 17846 68822 0.045% 0.395%
9 12070 13728 48196 0.038% 0.318%
10 13047 15031 53366 0.055% 0.452%
11 11933 14268 49990 0.039% 0.349%
12 10543 12371 43796 0.058% 0.598%
2019 1 11192 13056 45370 0.040% 0.335%
2 10940 12563 42594 0.038% 0.291%
3 13204 14906 54660 0.036% 0.327%
4 12754 14808 52608 0.043% 0.329%
5 13510 15486 53748 0.038% 0.348%
6 12755 13972 51938 0.042% 0.375%
7 10657 11824 40732 0.049% 0.407%
8 9938 11191 39042 0.047% 0.437%
9 10332 11711 41114 0.038% 0.315%
10 11544 12573 45260 0.031% 0.275%
11 12163 13089 46430 0.036% 0.328%
12 10834 12218 41796 0.040% 0.451%
2020 1 12432 14406 51486 0.063% 0.480%
2 12404 14033 50100 0.033% 0.280%
3 14105 16033 58402 0.038% 0.369%
4 11814 13644 46596 0.030% 0.271%
5 11658 13370 45256 0.047% 0.377%
6 13235 15100 52990 0.034% 0.292%
7 14963 17008 57558 0.026% 0.233%
8 13764 14974 53332 0.031% 0.277%
9 12427 13660 46116 0.025% 0.248%
10 6769 7649 25374 0.044% 0.441%
11 6177 6825 22626 0.052% 0.307%
12 12521 13572 47118 0.034% 0.328%